### v1

# AerialBooth: Mutual Information Guidance for Text Controlled Aerial View Synthesis from a Single Image

Abstract
We present a novel method, AerialBooth, for synthesizing the aerial view from a single input image using its text description. We leverage the pretrained text-to-2D image stable diffusion model as prior knowledge of the 3D world. The model is finetuned in two steps to optimize for the text embedding and the UNet that reconstruct the input image and its inverse perspective mapping respectively. The inverse perspective mapping creates variance within the text-image space of the diffusion model, while providing weak guidance for aerial view synthesis. At inference, we steer the contents of the generated image towards the input image using novel mutual information guidance that maximizes the information content between the probability distributions of the two images. We evaluate our approach on a wide spectrum of real and synthetic data, including natural scenes, indoor scenes, human action, etc. Through extensive experiments and ablation studies, we demonstrate the effectiveness of AerialBooth and also its generalizability to other text-controlled views. We also show that AerialBooth achieves the best viewpoint-fidelity trade-off though quantitative evaluation on $7$ metrics analyzing viewpoint and fidelity w.r.t. input image. Code and data is available at https://github.com/divyakraman/AerialBooth2023.
Figure 1: Given a single input image and its text description, AerialBooth generates an aerial view of the given image. It leverages pretrained text to 2D image models and does not require access to any paired data or explicit 3D or multi-view information during training or test time.
1 Introduction
Widely available 2D text-to-image models such as Stable Diffusion [49], trained on large-scale text-image data, contain immense knowledge of the 3D world. They are capable of generating various viewpoints of objects. Consequently, it can be inferred that text, along with 2D images, form a rich representation of the 3D space. However, given an in-the-wild image and its text description, the vanilla stable diffusion model may not be able to generate different viewpoints of the scene in the image due to limitations of the expressiveness of the text. Image personalization and editing approaches such as DreamBooth [50], Imagic [25] are successful in performing a wide range of non-rigid transformations including text-controlled view synthesis. However, the large translation required for novel view synthesis, especially aerial view, makes them sub-optimal [42] in generating the aerial viewpoint due to bias-variance tradeoff issues, especially when a single input image is provided. In this paper, we focus on text controlled aerial view synthesis from a single input image.
The usage of text to control the viewpoint makes the definition of ‘aerial view’ ambiguous since no camera parameters are explicitly involved. Without loss of generality, we consider any generated image that has a higher camera angle than the input image to be an aerial view of the image. The task of generating aerial views from a given image, also termed as cross-view synthesis, is an important computer vision task with applications [36] in cross-view localization and mapping [22], autonomous driving [7], augmented and virtual reality [15], 3D reconstruction [67], medical imaging [65], security and surveillance [2] and generation of realistic aerial view synthetic data for improved aerial view perception tasks. However, it is difficult to collect paired data and 3D metadata to train large generative models [42]. Consequentially, it is beneficial to develop an unsupervised method, that doesn’t use multi-view or depth information at train or test time, to generate aerial views from a single input image.
Prior work on text-controlled aerial view synthesis such as Aerial Diffusion [27] doesn’t work too well on complex scenes such as traffic scenes, indoor scenes and scenes with multiple objects and requires hyperparameter tuning for every input image. Moreover, many of the resulting images have residual diagonal artifacts. A closely related problem is novel view synthesis (NVS) and 3D synthesis [35, 42]. NVS methods sample different camera viewpoints by explicitly specifying the camera angle. However, they often need to trained on an enormous amount of large-scale datasets with scenes from different views and 3D data.
Main contributions.
We propose AerialBooth for aerial view synthesis from a single image. Our method leverages text-to-image diffusion models for prior knowledge and does not require any 3D or multi-view data. Moreover, our method performs test-time finetuning and hence does not require any additional training, it is easily generalizable to any in-the-wild image. The key novel components of our algorithm include:
  1. 1.
    Test-time finetuning: We first prepare the text-image embedding space by sequentially optimizing the text embedding and the diffusion model on the input image and its inverse perspective mapping in its close vicinity. Such a mechanism enables the model to learn the characteristics of the input image while acquiring adequate variance to generate aerial views.
  2. 2.
    Mutual Information Guided Inference: To generate the aerial view of the image using its text description, we propose a mutual information guidance formulation. Mutual information guidance steers the content in the generated aerial view towards the input image by maximizing the information content between the probability distributions of the input image and the generated aerial image, without imposing any pixel-to-pixel matching constraints between the two images with vastly different viewpoints.
Evaluation: To test our method, we collect a diverse set of synthetic images (from Stable Diffusion XL) and real images (from Unsplash), spanning from natural scenes, indoor scenes, to human actions and animations. Comparisons with prior work on single-image text-based view synthesis show that AerialBooth achieves a better viewpoint-fidelity (w.r.t. input image) trade-off to generate high-quality aerial images of the given input images, as demonstrated by our qualitative as well as quantitative results on $7$ metrics including CLIP Score [44], SSCD Score [40] and CLIP Directional Similarity [14]. We conduct extensive ablation experiments to show the usefulness of the inverse perpective mapping is useful in generating aerial views of images, and benefits of mutual information in guiding the contents in the generated image towards the contents of the input image. Our method can also be extended to generate more views that can be text-controlled (such as ‘side view’, bottom view’, ‘back view’), as evidenced by our results.
2 Related work
3D and Novel view synthesis.
Novel view synthesis [68, 64, 39] from a single image is an active area of research in generative AI. Many methods [60, 24, 17, 75] use NeRF based techniques. Nerdi [11] use language guidance with NeRFs for view synthesis. Many recent methods use diffusion [34, 57, 35, 43, 58, 55, 5, 52] to sample different views. 3D generation methods [42, 30, 71, 45, 8] use text to guide the reconstruction. All of these methods use large amounts of multi-view and 3D data for supervised training. Methods like Zero-1-to-3 [35] and Zero-123++ [55] use a pretrained stable diffusion [49] model, along with large data for supervised training, to learn different camera viewpoints. In contrast, we use the pretrained stable diffusion model, and don’t use any multi-view or 3D or any other kind of training data. Our method performs text-controlled view synthesis on a single input image, this paper doesn’t focus on explicitly controlling the (camera) angle of the generated view, the generated image is of arbitrary view, within the limits of the text guidance for viewpoint.
Image editing/ personalization.
Diffusion models have emerged as successful tools for single image editing and personalization. Methods such as DreamBooth [50], DreamBooth LoRA [21], HyperDreamBooth [51], Textual Inversion [16], Custom Diffusion [28] are able to generated personalized images of subjects. Image editing methods such as Imagic [25], Paint-by-Example [72], ControlNet [73], DiffEdit [9] are able to edit images to perform non-rigid transformations and also use exemplar signals for guidance. However, these methods can either generate aerial images with low fidelity w.r.t. the input image or generate high-fidelity images with viewpoints very close to the input image. In contrast, our method is able to execute larger view transformations than prior work, while preserving fidelity w.r.t. the input image.
Cross-view synthesis.
Prior work on cross-view synthesis [46, 61, 48, 63, 13, 36, 32, 56, 31, 47, 33, 69, 54, 1, 74] use paired data and modalities such as semantic maps, depth, multi-views, etc within their architectures. In this paper, we propose an unsupervised method for view synthesis using pretrained 2D text to image diffusion models. Aerial Diffusion [27] trains on homography images corresponding to Inverse Perspective Mapping (IPM) to generate aerial images by alternatingly sampling between prompts corresponding to viewpoint and the homography image. However, the diagonal artifacts in the resulting images indicate a need to utilize IPM better. Moreover, the quality of the generated images is very poor for complex scenes that typically contain more than one object. Our method is capable of generating high quality images for diverse complex scenes.
Guidance techniques in diffusion.
Guidance methods [19, 3, 12, 38] have been used to control and guide diffusion denoising towards semantic maps, image classes, etc. In this paper, we propose a novel diffusion guidance formulation to align the contents in two images that have vastly different views.
Figure 2: Overview. AerialBooth takes in a single input image $I_{S}$ to generate its aerial view using a pretrained text-to-image diffusion model, and without any 3D or multi-view supervision. It finetunes the diffusion model in two stages to optimize the text embedding and the diffusion model to reconstruct the input image and its inverse perspective mapping (or homography image) in close vicinity. Such a mechanism enables the incorporation of image specific knowledge within the model, while retaining its imaginative capabilities (or variance). At inference, AerialBooth uses mutual information guidance to maximize the information between the probability distributions of the generated image and $I_{S}$, to generate a high-fidelity image.
3 Method
We present AerialBooth  to generate aerial view images from a single input image $I_{S}$ and its text description $t_{S}$ (e.g. ‘a cosy living room’). The text description of the image can be obtained using the BLIP-2 model [29]. We leverage the pretrained text-to-2D image stable diffusion [49] model to serve as a strong prior. Furthermore, we do not use any training data or 3D/multi-view data. We present an overview of our method in Figure 2.
  • Finetuning the diffusion model: We perform a two-stage test-time finetuning on the input image $I_{S}$ to incorporate the knowledge pertaining to the input image within the pretrained model, while ensuring that the model does not get biased towards it. Specifically, we optimize the text embedding as well as the diffusion UNet sequentially on the input image and its inverse perspective mapping, in close vicinity.
  • Inference: To generate the aerial-view image, we use the target text description $t_{T}$ that takes the form ‘aerial view, ’ + $t_{S}$ (e.g. ‘aerial view, a cosy living room’). To ensure that the generated image has high fidelity with the input image, we use mutual information guidance.
Next, we describe our method in detail.
3.1 Test-time finetuning
The text-to-2D image stable diffusion model has knowledge of the 3D world as a consequence of the large amount of diverse data it has been trained on. It understands [53] different viewpoints, different styles, backgrounds, etc. Image editing and personalization methods such as DreamBooth [50], DreamBooth LoRA [21], Imagic [25], SVDiff [18] exploit this property to perform transformations such as making a standing dog sit and generating its image in front of the Eiffel tower. At a high level, the standard procedure adopted by these methods to generate edited or personalized images is to finetune the model, followed by inferencing. These methods are however not very successful in text-guided aerial view synthesis, which demands a large transformation. Specifically, directly finetuning the diffusion model on $e_{S}$ to reconstruct $I_{S}$ results in severe overfitting, where $e_{S}$ is the CLIP text embedding for $t_{S}$. This makes it difficult for the model to generate large variations to the scene required for aerial view synthesis.
We propose a two-step finetuning approach to enable the model to learn the characteristics of $I_{S}$, while ensuring sufficient variance for aerial view generation.
3.1.1 Optimized Text Embedding:
In the first step, we start from $e_{S}$ and compute the optimized text embedding $e_{opt}$ to reconstruct $I_{S}$ using a frozen diffusion model UNet using the denoising diffusion loss function $L$ [20].
$\min_{e_{opt}}\sum_{t=T}^{0}L(f(x_{t},t,e_{opt};\theta),I_{S}),$ (1)
where $t$ is the diffusion timestep and $x_{t}$ is the latents at time $t$. This formulation allows us to find the text embedding that characterizes $I_{S}$ better than the generic text embedding $e_{S}$. Next, to enable $e_{opt}$ accurately reconstruct $I_{S}$, we optimize the diffusion UNet using the denoising diffusion objective function. Note that we insert LoRA layers within the attention modules in the diffusion UNet and finetune only the LoRA layers with parameters $\theta_{LoRA}$, the rest of the UNet parameters are frozen,
$\min_{\theta_{LoRA}}\sum_{t=T}^{0}L(f(x_{t},t,e_{opt};\theta),I_{S}).$ (2)
While optimizing $e_{opt}$ instead of $e_{S}$ to reconstruct $I_{S}$ ensures lesser bias (or more variance), the embedding space is still not sufficiently conditioned to generate an aerial view of the image.
3.1.2 Inverse Perspective Mapping:
Inverse perspective mapping (IPM) [59] is a homography transformation from classical computer vision to generate the aerial-view of an image from its ground-view. Despite not being accurate, it can provide pseudo weak supervision for the generation of the aerial image and also add more variance to the embedding space. We denote the inverse perspective mapping of the input image by $I_{H}$, computed following [27]. In the second step of finetuning, to find the text embedding $e_{H}$ that best characterises $I_{H}$, we start from $e_{opt}$ and optimize the text embedding with a frozen diffusion model, similar to Equation 1. Finding $e_{H}$ in the vicinity of $e_{opt}$ instead of $e_{S}$ ensures that the text-image space corresponding to $e_{S}$ doesn’t get distorted to generate the poor quality IPM image. Next, we finetune the diffusion model using the denoising diffusion objective function to reconstruct $I_{H}$ at $e_{H}$, similar to Equation 2. Again, only the LoRA layers are finetuned, the rest of the UNet is frozen.
Note that we find $e_{opt}$ and $e_{H}$ by optimizing $e_{S}$ and $e_{opt}$, respectively for a small number of iterations. We need to ensure that $e_{S}$, $e_{opt}$ and $e_{H}$ are all in close vicinity. Our finetuning approach conditions the embedding space to encapsulate the details of $I_{S}$, while being able to generate large transformations required for the generation of the aerial image.
3.2 Inference
Our next step is to use the finetuned diffusion model to generate the aerial view image for the text prompt $t_{T}$. The text embedding for $t_{T}$ is $e_{T}$. Diffusion denoising, conditioned on $e_{T}$ is capable of generating aerial images corresponding to $I_{S}$. However, oftentimes, the contents of the generated aerial view image does not completely align with the contents of $I_{S}$. Consequently, to ensure high fidelity generations, our goal is to guide the contents of the aerial view image towards the contents of $I_{S}$.
Similarity measures such as L1 distance, cosine similarity are capable of providing this guidance. However, they are not invariant to viewpoint/ structure. Since we want the two images to be similar (while observed from different viewpoints), using metrics that impose matching at the pixel (or feature) level is not the best approach. Rather, it is judicious to use the probability distribution of the features.
In information theory, mutual information quantifies the ‘amount of information’ obtained about one random variable by observing the other random variable. Mutual information has been used [66, 37, 26, 70] to measure the similarity between images in various computer vision tasks such as medical image registration, frame sampling, etc. It yields smooth cost functions for optimization [62]. The mutual information between two probability distribution functions (pdf) $p(x),p(y)$ for two random variables $\mathcal{X},\mathcal{Y}$ is defined as $I(\mathcal{X},\mathcal{Y})=H(\mathcal{X})+H(\mathcal{Y})-H(\mathcal{X},\mathcal{Y})$ where $H(\mathcal{X}),H(\mathcal{Y})$ are the entropies of $p(x),p(y)$ and $H(\mathcal{X},\mathcal{Y})$ is the joint entropy. Entropy of a random variable $X$ is a measure of its uncertainity, $H(\mathcal{X})=-\sum_{x\in\mathcal{X}}p_{X}(x)log(p_{X}(x))$;
$H(\mathcal{X},\mathcal{Y})=-\sum_{(x,y)\in\mathcal{X,Y}}p_{XY}(x,y)log(p_{XY}(x,y)).$
Thus,
$I(\mathcal{X,Y})=-\sum_{(x,y)\in\mathcal{X,Y}}p_{XY}(x,y)\frac{log(p_{XY}(x,y))}{p_{X}(x)p_{Y}(y)}.$
Hence, mutual information, in some sense, measures the distance between the actual joint distribution between two probability distributions and the distribution under an assumption that the two variables are completely independent. Thus, it is a measure of dependence [23] and can be used to measure the information between two images. In order to maximize the similarity in content between $I_{S}$ and the generated aerial image, we maximize the mutual information between them. We define our mutual information guidance function as follows.
Let $z_{t}$ denote the predicted latents at timestep $t$. We denote $z_{0,t}$ as the latents of the final predicted image extrapolated from $z_{t}$ i.e. if the denoising were to proceed in a vanilla fashion in the same direction that computed $z_{t}$, the latents of the final predicted image would be $z_{0,t}$. At every step of sampling (except the final step), we wish to maximize the mutual information between $z_{0,t}$ and $z_{S}$ where $z_{S}$ are the latents corresponding to $I_{S}$. Hence, the guidance function we wish to maximize is, $G_{MI}=I(z_{0,t},z_{S}).$
The computation of mutual information requires us to compute the marginal and joint probability density functions (pdf) of $z_{0,t}$ and $z_{S}$. We construct 2D histograms of the latents (by reshaping the latents of size $C\times H\times W$ into $C\times HW$) and compute their marginal pdfs. The joint pdfs can then be computed from the marginal pdfs, which can be plugged into the formula for mutual information. Our next task is to use $G_{MI}$ to guide the generation of the aerial image.
Guidance techniques such as classifier-free guidance [19], universal guidance [3] and steered diffusion [38] modify the sampling method to guide the image generation with feedback from the guidance function. The gradient of the guidance function w.r.t. the predicted noise at timestep $t$ is an indicator of the additional noise that needs to be removed from the latents to steer the generated image towards the guidance signal. Synonymously, the gradient of the guidance function w.r.t. the predicted latents is an indicator of the direction in which the latents need to move in order to maximize their alignment with the guidance function. Specifically, at every step of sampling (except the final step), we modify the predicted latents $z_{t}$ as $\hat{z}_{t}=z_{t}-\lambda_{MI}\nabla_{z_{t}}(-1*G_{MI})$. Note that we use the negative of the mutual information to compute the gradients since we want to maximize the mutual information between the generated latents and the input image.
4 Experiments and Results
Figure 3: We show comparisons with the state-of-the-art image editing, personalization and text-controlled aerial view synthesis methods on the AerialBooth-Syn dataset. Our methods is able to generate images that are “more aerial”, while preserving the fidelity w.r.t. the input image.
Figure 4: We show comparisons with the state-of-the-art image editing, personalization and text-controlled aerial view synthesis methods on the AerialBooth-Real dataset. Our method achieves the best trade-off between generating images that are aerial and preserving the fidelity w.r.t. the input image.
Method CLIP A-CLIP SSCD DINO CLIP-I CLIPD A-CLIPD
Dataset: AerialBooth - Syn
DreamBooth LoRA $0.3069$ $0.1765$ $0.6260$ $0.6204$ $0.9058$ $0.0369$ $0.0334$
Imagic $0.3046$ $0.1748$ $0.5559$ $0.5846$ $0.9064$ $0.0315$ $0.0329$
Aerial Diffusion $0.2920$ $0.2082$ $0.2056$ $0.2946$ $0.751$ $0.0674$ $0.0905$
Ours ($\lambda_{MI}=1e-5$) $0.3107$ $0.1859$ $0.3921$ $0.4493$ $0.8636$ $0.0512$ $0.0486$
Ours ($\lambda_{MI}=1e-6$) $0.3108$ $0.1862$ $0.3928$ $0.4532$ $0.8678$ $0.0520$ $0.0469$
Dataset: AerialBooth - Real
DreamBooth LoRA $0.2929$ $0.1791$ $0.5501$ $0.5088$ $0.9000$ $0.0371$ $0.0332$
Imagic $0.2960$ $0.1773$ $0.5024$ $0.5048$ $0.9065$ $0.0340$ $0.0348$
Aerial Diffusion $0.2896$ $0.2084$ $0.1717$ $0.2614$ $0.7606$ $0.0711$ $0.0721$
Ours ($\lambda_{MI}=1e-5$) $0.3047$ $0.1887$ $0.3326$ $0.3884$ $0.8700$ $0.0547$ $0.0453$
Ours ($\lambda_{MI}=1e-6$) $0.3045$ $0.1871$ $0.3345$ $0.3915$ $0.8742$ $0.0564$ $0.0429$
Table 1: We compare AerialBooth with various state-of-the-art methods on quantitative metrics indicative of viewpoint, fidelity and directional similarity. Our method achieves the best viewpoint-fidelity trade-off amongst all prior work.
In this section, we present the results of our method along with extensive comparisons and ablations.
Data.
We collect a synthetic dataset, AerialBooth-Syn and a real dataset, AerialBooth-Real. Both datasets contain images across a wide variety of categories including indoor scenes, natural scenes, human actions, birds/ animals, animations, traffic scenes and architectures. AerialBooth-Syn contains $500$ images that were generated using Stable Diffusion XL [41]. To generate the text prompts for the generated images in AerialBooth-Syn, we used Large Languuage Models (LLMs) such as ChatGPT and Bard. AerialBooth-Real contains $139$ images downloaded from the Unsplash website, the text descriptions for these images were obtained using the BLIP-2 [29] model.
Training details.
We use the stable diffusion 2.1 model in all our experiments, ablations and comparisons. All our images (except for images in AerialBooth-Real) are at a resolution of $512\times 512$. With respect to $I_{S}$, we train the text embedding and the diffusion model for $1,000$ and $500$ iterations respectively, at a learning rate of $31e-$ 3 and $42e-$ 4 respectively. Using $1000$ iterations to optimize the text embedding ensures that the text embedding $e_{opt}$ at which $I_{S}$ is reconstructed is not too close to $e_{S}$, which would make it biased towards $I_{S}$ otherwise. Similarly, it is not too far from $e_{S}$ either, hence the text embedding space learns the characteristics of $I_{S}$. With respect to $I_{H}$, we train the text embedding and the diffusion UNet for $500$ and $250$ iterations respectively. We want $e_{H}$ to be in close vicinity of $e_{opt}$; we train the diffusion model for just $250$ iterations so that the model does not completely overfit to $I_{H}$. The role of $I_{H}$ is to create variance and provide pseudo supervision, it is not an accurate approximation of the aerial view. We set the hyperparameter for mutual information guidance at $51e-$ 5 or $61e-$ 6, the inference is run for $50$ steps.
Quantitative evaluation metrics.
No concrete evaluation metrics exist for this task. We follow prior work on text-based image editing and personalization to evaluate our method on the following $7$ metrics:
  • Viewpoint and text alignment: We use the text description ‘aerial view, ’ + $t_{S}$ and ‘aerial view’, along with the generated image, to compute the CLIP-Score [44] and the A-CLIP Score respectively. The former indicates alignment of the generated image with detailed textual description of the image; the latter focuses more on the viewpoint.
  • Fidelity: To evaluate the alignment of the contents of the generated aerial-view image with the input image, we compute the CLIP-I score [50] which measures the cosine similarity between the embeddings of the aerial-view image and the input image in the CLIP space. We also use the self-supervised similarity detection metrics, DINO [6] and SSCD [40].
  • CLIP Directional Similarity: CLIP directional similarity [14] measures the consistency of the change between two images, $I_{S}$ and $I_{T}$, in the CLIP space with the change between the two image captions (dictating the transformation from $I_{S}$ to $I_{T}$). We compute two versions of this score, CLIPD score and the A-CLIPD score. CLIPD score uses ‘aerial view, ’ + text as the target text, and ‘A-CLIPD score’ uses ‘aerial view’ as the target text.
Higher values are desired for each of these metrics. Viewpoint faithfulness and fidelity w.r.t input image are a direct result of the bias-variance trade-off of the model. High values for both are desired. However, as noted by Blau et. al. [4], maximizing both is not straightforward; inevitably one of the factors will degrade in response to the improvement in the other. We generate $5$ aerial images, with random noise initializations, and choose the image with the highest CLIP + SSCD score (since CLIP is an indicator of viewpoint alignment and SSCD score measures the fidelity w.r.t. the input image).
Figure 5: We conduct ablation experiments to show the effectiveness of Inverse Perspective Mapping (IPM) and the mutual information guidance. IPM helps in the generation of images that are aerial, and mutual information guidance better preserves the contents of the input image.
State-of-the-art comparisons.
We compare our method with DreamBooth LoRA [50], a text-based image personalization method; Imagic LoRA [25], a text-based image editing method; and Aerial Diffusion LoRA [27], a method for text-based aerial image generation from a single image. We keep the backbone stable diffusion model, image prompts, training details, and evaluation method consistent across all comparisons.
4.1 Results
In this subsection, we analyze the results of our method and prior art on AerialBooth-Syn and AerialBooth-Real.
We show qualitative results on AerialBooth-Syn in the first row of Figure 1 and in Figure 3. We show qualitative results on AerialBooth-Real in the second row of Figure 1 and in Figure 4. Our method is able to generate aerial views of the input images across a diverse set of scenes. Our method generates results that are more aerial in viewpoint than DreamBooth and Imagic, while being consistent with the contents of the input image. Aerial Diffusion is unable to generate good quality images of scenes that have many objects. Our method is able to deal with complex scenes as well as modify the viewpoint.
We show the quantitative results in Table 1. Our method achieves a higher CLIP score than all prior work, indicating that it is able to generate an aerial view of the scene with contents dictated by the text better than prior work. The A-CLIP score achieved by our method is higher than that of DreamBooth and Imagic, indicating better conformance to the aerial viewpoint. Even though the A-CLIP score of AerialBooth is lower than than of Aerial Diffusion, Aerial Diffusion generates poor quality images for scenes with more than one object and also has diagonal artifacts in its generated images, as we observe from the other metrics and qualitative results, thus, offsetting its high A-CLIP Score.
CLIP-I and self-supervised metrics such as SSCD and DINO are not viewpoint invariant. SSCD is a better metric [40] than DINO. In many cases, since Imagic and DreamBooth generate views close to the input view, instead of an aerial views, it is natural for them to have higher SSCD and DINO Score. Our method has a much higher SSCD and DINO score than Aerial Diffusion, showing considerable improvement over prior work in retaining the fidelity w.r.t. the input image, while modifying the viewpoint.
CLIPD and A-CLIPD measure the directional change in the generated image w.r.t. the text prompt describing the transformation. Higher values indicate the image moving more towards the aerial viewpoint. Our method achieves higher CLIP scores than Imagic and DreamBooth, indicating its success in generating aerial images. Synonymous to our analysis for CLIP and A-CLIP, while the CLIPD and A-CLIPD scores are higher for Aerial Diffusion than our method, Aerial Diffusion is unable to generate good quality images as evidenced by the CLIP-I score.
In summary, AerialBooth achieves the best viewpoint-fidelity trade-off amongst all prior work.
4.2 Ablation studies
We show ablation experiments in Figure 5. In the second column, we show results of our model where it is neither finetuned on the homography image nor uses mutual information guidance for sampling. Thus, the text embedding for the input image, followed by the diffusion UNet are finetuned and the diffusion model generates the aerial image by diffusion denoising, without any mutual information guidance. Many of the generated images either have low fidelity or have low correspondence to the aerial viewpoint. In the experiment in the third column, we add mutual information guidance to the model in column 2. We see higher fidelity (than column 2) of the generated images w.r.t. the input image. In the fourth column, we add the homography image finetuning step to the model in column 2, but do not use mutual information guidance at inference. The generated images, in many cases, are aerial, but have lower fidelity w.r.t. the input image. In the final column, we show results with our full model. The generated images achieve the best trade-off between the viewpoint being aerial and fidelity w.r.t. the input image, in comparison to all ablation experiments. We report a table with all quantitative metrics for the ablation experiments on AerialBooth-Syn and AerialBooth-Real in the supplementary pdf. A summary of the analysis is as follows:
  • Effect of $I_{H}$: To study the effect of $I_{H}$, we compare the A-CLIP, CLIPD, A-CLIPD scores for the full model vs model w/o $I_{H}$ and model w/o $I_{H}$ w/o $G_{MI}$ vs model w/o $G_{MI}$. The scores in all cases where $I_{H}$ is not present are lower. For instance, on AerialBooth-Real, the CLIP-D scores for the full model and model w/o $I_{H}$ are $0.0564$ and $0.0444$ respectively. Similarly, the CLIP-D scores for model w/o $I_{H}$ w/o $G_{MI}$ vs model w/o $G_{MI}$ are $0.0550$ and $0.0445$ respectively.
  • Effect of $G_{MI}$: To study the effect of $I_{H}$, we compare the SSCD, DINO, CLIPD, A-CLIPD scores for the full model vs model w/o $G_{MI}$ and model w/o $I_{H}$ w/o $G_{MI}$ vs model w/o $I_{H}$. The scores in all cases where $G_{MI}$ is not present are lower. For instance, on AerialBooth-Real, the SSCD scores for the full model and model w/o $G_{MI}$ are $0.3345$ and $0.3204$ respectively. Similarly, the CLIP-D scores for model w/o $I_{H}$ w/o $G_{MI}$ vs model w/o $I_{H}$ are $0.4013$ and $0.4066$ respectively.
Figure 6: AerialBooth can be extended to generate other text-controlled views as well.
Figure 7: We compare with latest related work on novel view synthesis: Zero-1-to-3[35] and Zero123++ [55]. Both methods use the pretrained stable diffusion model and the 3D objects dataset, Objaverse [10] for training. Our method uses just the pretrained stable diffusion model to generate better results for the task of aerial view synthesis from a single image.
4.3 Text controlled view synthesis: Other views
AerialBooth can be extended to generate other text-controlled views such as side view, bottom view and back view. For the results shown in Figure 19, we modify the target text $t_{T}$ to indicate different viewpoints, and retain the other finetuning/ inference details.
4.4 NVS Comparisons
We compare with latest work on stable diffusion based novel view synthesis from a single image in Figure 7. Zero-1-to-3[35] and Zero123++ [55], both, train on large amounts of multi-view and 3D data from Objaverse [10] contain $800K+$ 3D models; in addition to leveraging a pretrained stable diffusion model. In contrast, our method does not use any multi-view or 3D data for training and performs testtime finetuning to generate better results. However, while it is possible to explicitly control the camera angle and generate 3D objects in Zero-1-to-3[35] and Zero123++ [55], our method generates 2D images from arbitrary camera angles, within the realms of the text description of the viewpoint and consistent with the contents of the input image.
4.5 Comparison with other metrics for guidance.
We compare with two other metrics for diffusion guidance at inference: (i) L2 distance between the features of the generated image and the input image, (ii) a metric inspired by Wasserstein distance or Earth Mover’s distance, for which we compute the distance between the histograms of the probability distributions of the two images. Our mutual information guidance method is better at preserving the fidelity w.r.t the input image, as evidenced by higher SSCD scores. The SSCD score on AerialBooth-Real for Wasserstein guidance, $L2$ guidance, and mutual information guidance are $0.3181$, $0.3224$ and $0.3345$ respectively.
More results.
Please refer to the supplementary material for more analysis.
5 Conclusions, Limitations and Future
We present a novel method for aerial view synthesis. Our goal is to leverage pretrained text-to-image models to advance the frontiers of text-based view synthesis from a single image without any additional 3D or multi-view data at train/test time. We define aerial view to be any view that has a higher camera angle than the input image. It will be interesting to incorporate explicit control over the camera angle, along with text-controlled view synthesis, in a data efficient manner. Methods like Zero123++ [55] and Zero NVS [52], after being trained on large-scale 3D data, have some amount of generalization capabalities. These pretrained models can be combined with our method to control camera views for novel view synthesis. The inherent bias-variance tradeoff results in small changes in some generated aerial image results, further improving the fidelity w.r.t. input image and the viewpoint is a direction for future work. Inverse perspective mapping serves as an important tool for aerial view generation – working on similar homography projections to synthesize other views can be explored as well. Our mutual information guidance formulation can be used in other problems pertaining to image editing and personalization to maximize the information in the probability content between two images. Using generated data for various applications such as cross-view mapping, 3D reconstruction, data augmentation for training autonomous systems is also an avenue for future work.